OpenStudy (beststudent):

Which of the following groups of side lengths would form a right triangle? A. 4 cm, 5 cm, 6 cm B. 7 cm, 24 cm, 25 cm C. 8 cm, 16 cm, 30 cm D. 10 cm, 40 cm, 50 cm

OpenStudy (beststudent):

hey

OpenStudy (anonymous):

so you have to use the pythgorean theorem which is a^2+b^2=c^2

OpenStudy (beststudent):

yes

OpenStudy (anonymous):

ok so its just a guess and check and c must always be the biggest side

OpenStudy (beststudent):

so it c

OpenStudy (anonymous):

also the 7 24 25 is a special right triangle so you know that will always work along with the 3 4 5 triangle and others

OpenStudy (anonymous):

so for e first one it would be 4^2+5^2=6^2 which is 16+25=36 ------> 41 doesn't equal 36 therefore its not that one

OpenStudy (beststudent):

A right triangle has sides measuring 6 units, 8 units, and 10 units. Which of the following shows the Pythagorean relationship of the areas of the squares drawn along each side? A. 12 + 16 = 28 B. 18 + 24 = 42 C. 24 + 76 = 100 D. 36 + 64 = 100

OpenStudy (anonymous):

no its not c because 8^2+16^2 doesn't equal 30^2

OpenStudy (anonymous):

the answer for the previous one is b

OpenStudy (anonymous):

the new ques is 6^2+8^2=10^2 so you just have to compute that and you will find the answer